An experienced accountant can prepare a tax return in 11 hours. A novice accountant can do the job in 17 hours. How long will it take them to do the job working together? answer in fraction an hr

Answer by Boreal(15235)

Estimate 9% for first and 6% for second, so answer should be about 15% per hour or between 6 and 7 hours.
(1/11) is done per hour by experienced. (1/17) by novice.
that equals 1.
(x/11)+(x/17)=1
multiply everything by 187, common denominator
17x+11x=187
28x=187
x=187/28 hours, which is between 6 and 7.